OpenStudy (anonymous):

if a set of markers is placed in rows of 4 each,there are two markers left over, if in rows of 5 each , there are 3 left over;and if in row of 7 there are 5 lefft over.what is the smallest number of markers that the set could contain?

OpenStudy (littlebird):

This is what I found on the Internet. Hopefully it is correct. :)

OpenStudy (littlebird):

@drsuganyah look at the picture!

OpenStudy (anonymous):

138 markers is my answe

OpenStudy (littlebird):

Awesome

OpenStudy (anonymous):

the lcm of 4 , 5, 7 is 140 the difference between the asked (4 remainder 2 so 4-2 =2, 5-3=2, 7-5=2, so the difference is 2 in common , i subtracted 140-2 to get 138.

OpenStudy (anonymous):

thanks little bird

OpenStudy (littlebird):

What math class do you learn this in? I honestly had no idea how to do this until I googled it.

OpenStudy (anonymous):

actually i take math tuition, this was a question sent for a 7th grade math student PreAlgebra level....charlotte private school

OpenStudy (littlebird):

I never had to do this for algebra, or prealgebra

OpenStudy (anonymous):

which grade are you/

OpenStudy (littlebird):

I'm now taking calculus.

OpenStudy (anonymous):

good... keep going....

OpenStudy (anonymous):

good...proud

OpenStudy (anonymous):

All the Best.Take care. see you Bye